State the number of atoms of each element on each side of the following equation. CH4(g)+2O2(g)→CO2(g)+2H2O(g)
1
Expert's answer
2017-07-31T11:53:06-0400
In the following chemical reaction:
CH4(g)+2O2(g)→CO2(g)+2H2O(g)
On the left side the number of each element:
Carbon C - 1 atom
Hydrogen H – 4 atoms
Oxygen O – 2*2 = 4 atoms
On the rigth side the number of each element:
Carbon C - 1 atom
Hydrogen H – 2*2 = 4 atoms
Oxygen O – 2 + 2*1 = 4 atoms
It means that there is equilibrium state.
Answer: C – 1; H – 4; O - 4.